OpenStudy (anonymous):

How would I graph this function? Y=(x+3)^2-4

Vocaloid (vocaloid):

well, the ^2 tells us that this is a parabola, and since the leading coefficient is positive, we have an upwards facing parabola

Vocaloid (vocaloid):

so, let's start off with the parent function y = x^2, which looks like this:

Vocaloid (vocaloid):

now, the graph y = (x+3)^2 - 4 is just a variation y = x^2. the +3 inside the parentheses tells us to shift the graph left 3 units, and the -4 tells us to shift the graph down 4 units